# Connecting the Dots: Evaluating Abstract Reasoning Capabilities of LLMs Using the New York Times Connections Word Game

### v5

@@ 4.1 LLMs as Game Players
To test the capabilities of large language models in solving the Connections game, we rely on recent advancements in in-context learning and chain-of-thought prompting Wei et al. (2022). We provide 3 complete examples in our few-shot prompt along with rules and common strategies that players must use to solve the game. We also elicit chain-of-thought reasoning Wei et al. (2022) requiring models to explain their groupings and categories chosen. Formulation of the prompt involved trial and error; the first iteration of the prompt included the Connections game instructions provided by the New York Times (Liu, 2023a), and included three demonstrations with gold labels asking the LLM to explain its reasoning in a step-by-step manner Wei et al. (2022). We ran this first prompt with a few games on a development set of 30 games (different from our test set), using the 4 LLMs. After identifying commonalities in the types of errors made by the LLMs while playing the game, we added additional instructions about the game, specified the response format, and included some tips from a NYT article about playing Connections (Aronow and Levine, 2023). The entire prompt is in Appendix A. To ensure consistency and fairness in performance, we prompt 4 LLMs — Gemini 1.5 Pro, Claude 3 Opus, GPT-4o, and Llama 3 70B —with the same input and use the default sampling parameters (temperature and top_p). We use the scoring schema outlined in Section 4.3 to evaluate how all models perform in solving 200 Connections games spanning from June 15, 2023 to January 1, 2024.
@@ 4.2 Humans as Game Players
Alongside LLMs, we recruited 17 human evaluators in two subgroups: 12 novice players with little to no prior experience playing Connections and 5 expert or regular Connections players. The novice and the expert evaluators were peers of the authors of the paper who volunteered to participate without any payment.
We designed a human evaluation interface and randomly sampled 100 games from our test set. Appendix E has more information about the interface. The first screen displays a shortened version of the instructions from the LLM’s prompt so as to not overwhelm the human players. To ensure that the humans solve the game in a manner comparable to the LLMs setup, they were given one try to fully solve the game (i.e., make all 4 categorizations). This aligns with Todd et al. (2024)’s challenge mode.
Playing these games is a significant cognitive burden. As such, each novice human evaluator played around 8-12 distinct games for a total of 100 randomly sampled games out of the 200 games in the test set, and expert participants each played 10 games for a total of 50 games from the subset of 100 games played by novices.
@@ 4.3 Evaluation Criteria
Our scoring schema was developed as a means to numerically interpret the outcome of each Connections game and standardize comparison across LLMs and human players. We outline two processes to obtain clustering and categorical reasoning scores of a game of Connections.
@@ 4.3.1 Clustering Score
The clustering score evaluates the ability to correctly group together all the words in the game. We consider two clustering scores. The first, or the unweighted clustering score, is calculated independently of the categories’ supposed difficulty. In this simple scoring mechanism, we allocate one point for each correct cluster (when all 4 words in the group classified by the LLM or human correspond to the 4 words in the gold category/grouping). Ideally, a player’s score should be close to the maximum of 4, signifying that all 4 groups were correctly identified. The equation is as follows:
@@ 4.3.2 Categorical Reasoning
While the weighted and unweighted clustering scores are calculated for LLMs and humans, the categorical reasoning score is used only for the LLMs’ responses. If all 4 words in a category are correctly identified by an LLM, we conduct further analysis to evaluate whether the LLM reasoned correctly why the words in the groups belong together. We make this distinction in our evaluation so that — in conjunction with the taxonomy knowledge for Connections categories (Section 3.2) — we can assess the types of reasoning that the LLMs are most or least adept in. Since our prompt asks the LLM to include the category name and share the reasons why it grouped words together, we can evaluate whether the LLM’s reasoning in its response is semantically analogous to the gold NYT Connections-provided category name. The decision of semantic equivalence between LLMs output and gold is done manually by a human judge to ensure accuracy.
@@ 5.1 LLM performance
Overall, we find that GPT-4o performs best across all 200 games. Figure 2 shows the unweighted clustering scores for all 4 LLMs. GPT-4o has the lowest percentage of games in which it made no correct clustering (53 out of 200) and the most games solved perfectly (16 out of 200). Claude 3 Opus is a close second for perfectly solved games at 13 out of 200. Gemini 1.5 Pro performs the worst overall. While it could not make any correct clusters for half of the games, it was able to solve 9 games perfectly, outperforming Llama 3 70B’s 6 games solved perfectly.
In terms of weighted clustering scores for each model (Figure 3), Gemini 1.5 Pro and Llama 3 70B show similar results. Most of their scores are concentrated before 2, showing that these models showed a higher ability to correctly classify the easiest or second easiest categories. GPT-4o and Claude 3 Opus also showed similar results, with most of their weighted clustering scores concentrated before 4, meaning they were better at classifying more and harder categories. Weighted clustering scores are very rarely represented in all the models. Appendix D.2 contains a more detailed breakdown.
@@ 5.2 Human Performance
In human performance, we measure both novice and expert players against the best overall performing GPT-4o. For the 100 games played by novices and 50 games played by experts, we compare the same 100 and 50 games played by GPT-4o.
@@ 5.2.1 Novice Players
In the 100 games that the novice players completed, their average unweighted clustering score was 1.38, marginally better than GPT-4o’s average of 1.17 on the same 100 games. GPT-4o and novice humans also had similar weighted clustering score distributions. More details are in Appendix D.1. Due to the setup of the human interface, humans could not receive a clustering score of 3 (if humans correctly solve 3 groupings, the 4th is also correct). Because of GPT-4o’s imperfect instruction-following abilities (repeating or omitting a word), it was still able to obtain a clustering score of 3, as shown in Figure 4.
@@ 5.2.2 Expert Players
Expert human players performed significantly better than novices and GPT-4o, with an average clustering score of 3 compared to GPT-4o’s 1.22 (on the same 50 games) and an average weighted clustering score of 7.4 compared to GPT-4o’s 2.32. The distribution of weighted clustering scores is also far more right-skewed (see Appendix D.1 for more). Figure 5 shows that experts perfectly solve over 60% of the 50 games, while GPT-4o only fully solved 5% of the games.
@@ 6 Discussion
6.1 What type of reasoning is hardest for LLMs and humans?
To answer this question we rely on our taxonomy of reasoning types we introduced in Section 3.2. The breakdowns of the reasoning types for the 800 categories in our 200-game dataset are shown in Table 2. The patterns in performance across the types of reasoning parallel the LLMs’ overall performance for the most part, with Claude 3 Opus’s performance in Multiword Expressions defying this pattern. The performance in reasoning categories across models is ranked from best to worst as follows: . The model’s performance in these categories corresponds to the frequency with which each category appears, except the category of Multiword Expressions. Only Claude 3 Opus has a greater than zero success rate (4 out of 77) in Multiword Expressions. In Connections games with Multiword Expressions, they are usually the purple category (most difficult), while semantic and associative reasoning appear most often as yellow or green (easier) categories. For Combined categories, no models output correct clusters.
We find that both novice and expert human players are better at all types of reasoning compared to GPT-4o, although neither cluster any Combined Knowledge categories successfully. Novice performance in reasoning categories across models is ranked from best to worst as follows: . The greatest difference between GPT-4o and novices is in Multiword Expressions (difference of around 20%) and Linguistic Knowledge (difference of 25%). Expert performance ranked from best to worst: .
For both LLMs and humans, the Combined Knowledge category seems hardest to grasp. Though experts rank low in Multiword Expressions compared with other types of reasoning, they still perform quite highly, with over 60% of Multiword Expression categories grouped correctly. However, perhaps because of a lack of familiarity with the game and types of categories, novice players, like LLMs, struggle with Multiword Expression, achieving an accuracy of just over 20%. Further breakdowns are in Appendix D.
6.2 How much do distractors prevent both LLMs and humans from correct categorization?
The Connections game is often formulated with item overlap in mind, according to the Connections puzzle creator (Liu, 2023b). These distractors, or red herrings, make the game far more challenging. Red herrings can appear in two ways — as a red herring category or red herring word. In the former instance, 3 ultimately unconnected words seem to form a category of their own with 1 word missing. In the latter, a category seems applicable to more than 4 words, but the extras belong to a separate grouping. Examples of each of these types of red herrings are in Appendix C.
@@ 6.3 How often do LLMs group the words correctly but present incorrect reasons?
Gemini 1.5 Pro | 0.78 |
Claude 3 Opus | 0.87 |
GPT-4o | 0.86 |
Llama | 0.76 |
The average ratios in Table 3 are fairly high, close to or above 80%. The highest overall performing models GPT-4o and Claude 3 Opus have the highest ratios as well. Though it is fairly uncommon that a model will correctly group without correctly reasoning, there are very few instances where models received both a clustering score of 4 (fully solved game) and a categorical reasoning score of 4.
@@ 6.4 How can future work improve on such a benchmark?
Certain strategies grounded in reflective thinking could improve performance on such a benchmark. Instead of greedily choosing the first grouping, identifying red herring words or categories first could prevent the possibility of misclassification. Allowing LLMs to solve the game one category at a time and incorporating the feedback present to humans in the NYT Connections game, including whether a grouping is correct (and what difficulty level it is by color), incorrect, or one word away from a correct grouping, may improve performance as well. Retrieval Augmentation from WordNet or dictionaries for lexical connotations Allaway and McKeown (2020) could further improve such categorization. Finally, creating synthetic training data and training an LLM on this task could further close the gap between expert human and LLM performance. We leave such exploration for future work.
@@ 7 Conclusions
We introduced NYT Connection games as a benchmark to test abstract reasoning in state-of-the-art LLMs and evaluate their performance against expert and novice human players. We find that GPT-4o performs best, although it is still no match for expert human players. By examining the performance through our knowledge taxonomy, we obtain a more solid understanding of areas in which LLMs can improve to solve clustering scores. They are fairly deficient in certain types of reasoning required to be a skilled Connections player. Although most possess adequate semantic and associative reasoning capabilities, they struggle with Multiword Expressions and Combined Knowledge categories. Additional struggles arise because they cannot identify red herrings and use step-by-step logic to work around them. Ultimately, we find that excelling in Connections means having a breadth of different knowledge types, and LLMs are unfortunately not yet suited for the task.
@@ 8 Limitations
Many of the limitations in this section stem from the lack of data available for Connections games and disparities in the comparison between LLMs and humans. Because it is a fairly recent invention and only one puzzle is released per day, there are only a few hundred games available. Since there are some category patterns learned through frequent play, ideally, a model trained on past Connections games might bridge the gap between LLM are expert human evaluators performance.
We acknowledge that human evaluators were not required to add justifications for the groupings they made. This could have made performance comparisons between humans and LLMs for the types of reasoning more equal. Additionally, because a score of 3 was impossible in the human evaluation interface, we cannot be certain that humans were adept in the type of knowledge of their last category grouped, as this could simply been a matter of grouping all options left. Other limitations of human evaluators include that because they were all peers or acquaintances of the paper’s authors, sampling bias could exist. Though the age range of the humans recruited was 14-60, other demographic factors that may not have been accounted for in this sample.
@@ Appendix A Prompt
Categories will always be more specific than ’5-LETTER-WORDS’, ‘NAMES’, or ’VERBS.’
@@
Words: [’DART’, ‘HEM’, ‘PLEAT’, ‘SEAM’, ‘CAN’, ‘CURE’, ‘DRY’, ‘FREEZE’, ‘BITE’, ‘EDGE’, ‘PUNCH’, ‘SPICE’, ‘CONDO’, ‘HAW’, ‘HERO’, ‘LOO’]
@@
Ways to preserve food: [CAN’, ‘CURE’, ‘DRY’, ‘FREEZE’]
@@
Sharp quality: [’BITE’, ‘EDGE’, ‘PUNCH’, ‘SPICE’]
@@
Birds minus last letter: [’CONDO’, ‘HAW’, ‘HERO’, ‘LOO’]
@@
Words: [’COLLECTIVE’, ‘COMMON’, ‘JOINT’, ‘MUTUAL’, ‘CLEAR’, ‘DRAIN’, ‘EMPTY’, ‘FLUSH’, ‘CIGARETTE’, ‘PENCIL’, ‘TICKET’, ‘TOE’, ‘AMERICAN’, ‘FEVER’, ‘LUCID’, ‘PIPE’]
@@
Shared: [’COLLECTIVE’, ‘COMMON’, ‘JOINT’, ‘MUTUAL’]
@@
Rid of contents: [’CLEAR’, ‘DRAIN’, ‘EMPTY’, ‘FLUSH’]
@@
Associated with “stub”: [’CIGARETTE’, ‘PENCIL’, ‘TICKET’, ‘TOE’]
@@
__ Dream: [ ’AMERICAN’, ‘FEVER’, ‘LUCID’, ‘PIPE’])
@@
Words: [’HANGAR’, ‘RUNWAY’, ‘TARMAC’, ‘TERMINAL’, ‘ACTION’, ‘CLAIM’, ‘COMPLAINT’, ‘LAWSUIT’, ‘BEANBAG’, ‘CLUB’, ‘RING’, ‘TORCH’, ‘FOXGLOVE’, ‘GUMSHOE’, ‘TURNCOAT’, ‘WINDSOCK’]
@@
Parts of an airport: [’HANGAR’, ‘RUNWAY’, ‘TARMAC’, ‘TERMINAL’]
@@
Things a juggler juggles: [’BEANBAG’, ‘CLUB’, ‘RING’, ‘TORCH’]
@@
Words ending in clothing: [’FOXGLOVE’, ‘GUMSHOE’, ‘TURNCOAT’, ‘WINDSOCK’]

### v6

@@ 4.1 LLMs as Game Players
To test the capabilities of large language models in solving the NYT Connections game, we rely on recent advancements in in-context learning and chain-of-thought prompting Wei et al. (2022). We provide 3 complete examples in our few-shot prompt along with rules and common strategies that players must use to solve the game. We also elicit chain-of-thought reasoning Wei et al. (2022), requiring models to explain their groupings and categories chosen. Formulation of the prompt involved trial and error; the first iteration of the prompt included the NYT Connections game instructions provided by the New York Times (Liu, 2023a), and included three demonstrations with gold labels asking the LLM to explain its reasoning in a step-by-step manner Wei et al. (2022). We ran this first prompt on a development set of 30 games, using 5 LLMs. After identifying commonalities in the types of errors made by the LLMs while playing the game, we added additional instructions, specified the response format, and included some tips from a NYT article about playing the NYT Connections game (Aronow and Levine, 2023). The entire prompt is in Appendix A. To ensure consistency and fairness in performance, we prompt 5 LLMs — Gemini 1.5 Pro, Claude 3.5 Sonnet, GPT-4o, and Llama 3.1 405B and Mistral Large 2 — with the same input. We use the default sampling parameters (temperature and top_p) and the scoring schema outlined in Section 4.3 to evaluate how all models perform in solving 438 NYT Connections games spanning from June 2023 to August 2024.
@@ 4.2 Humans as Game Players
Alongside LLMs, we recruited 17 human evaluators in two subgroups: 12 novice players with little to no prior experience playing NYT Connections and 5 expert or regular NYT Connections players. The novice and the expert evaluators were peers of the first four authors, who volunteered to participate without any payment. We designed a human evaluation interface and randomly sampled 100 games from our test set. Appendix E has more information about the interface. The first screen displays an abridged version of the instructions from the LLM’s prompt so as to not overwhelm the human players. To ensure that the humans solve the game in a manner comparable to the LLMs setup, they were given one try to solve the game (i.e., make all 4 categorizations at once). This aligns with Todd et al. (2024)’s challenge mode.
Playing these games is a significant cognitive burden. As such, each novice human evaluator played around 8-12 distinct games for a total of 100 randomly sampled games out of the 438 in the test set, and expert participants each played 10 games for a total of 50 randomly sampled games.
@@ 4.3 Evaluation Criteria
Our scoring schema was developed as a means to numerically interpret the outcome of each NYT Connections game and to standardize the comparison across LLMs and human players. We outline two processes to obtain clustering and categorical reasoning scores for a game of Connections.
@@ 4.3.1 Clustering Score
The clustering score evaluates the ability to correctly group all the words in the game. We consider two clustering scores. The first, or the unweighted clustering score, is calculated independently of the categories’ supposed difficulty. In this simple scoring mechanism, we allocate one point for each correct cluster (when all 4 words in the group classified by the player match the 4 words in the gold category/grouping). Ideally, a player’s score should be close to the maximum of 4, signifying that all 4 groups were correctly identified. The equation is as follows:
@@ 4.3.2 Categorical Reasoning
While the weighted and unweighted clustering scores are calculated for LLMs and humans, the categorical reasoning score is used only for the LLMs’ responses. If all 4 words in a category are correctly identified by an LLM, we conduct further analysis to evaluate whether the LLM reasoned correctly why the words in the groups belong together. We make this distinction in our evaluation so that—in conjunction with the taxonomy of knowledge for NYT Connections categories (Section 3.2)—we can assess the types of reasoning that the LLMs are most or least adept in. Since our prompt asks the LLM to include the category name and share the reasons why it grouped words, we can evaluate whether the reasoning in its response is semantically analogous to the gold NYT Connections-provided category name. The decision of semantic equivalence between LLMs output and gold is done manually by a human judge to ensure accuracy.
@@ 5.1 LLM performance
Overall, we find that Claude 3.5 Sonnet performs best across thw 438 games. Figure 3 shows the unweighted clustering scores for all 5 LLMs. Claude has the lowest percentage of games in which it made no correct clustering, at about 20%, and most games solved perfectly at 18%. Mistral Large 2 performs the worst overall. It could not make any correct clusters for 42% of the games, and it perfectly solved the least amount of games.
In terms of weighted clustering scores for each model (Figure 4), Gemini 1.5 Pro and Mistral Large 2 show similar results. Most of their scores are concentrated before 2, showing that these models had a higher ability to correctly classify the easiest or second easiest categories. GPT-4o and Claude 3.5 Sonnet had most of their weighted clustering scores concentrated before 5, meaning they were better at classifying more and harder categories. Weighted clustering scores 8 are very rarely represented in all the models. Appendix D.2 contains a more detailed breakdown.
@@ 5.2 Human Performance
In human performance, we measure both novice and expert players against the best overall performing Claude 3.5 Sonnet. For the 100 games played by novices and 50 games played by experts, we compare the same games played by Claude 3.5 Sonnet.
@@ 5.2.1 Novice Players
In the 100 games that the novice players completed, their average unweighted clustering score was 1.37, slightly worse than Claude’s average of 1.52 in the same 100 games. Claude and novice humans also had similar weighted clustering score distributions. More details are in Appendix D.1. Due to the setup of the human interface, humans could not receive a clustering score of 3 (if humans correctly solve 3 groupings, the fourth is also correct). Because of Claude’s imperfect instruction-following abilities (repeating or omitting words), it was still able to obtain a clustering score of 3, as shown in Figure 5.
@@ 5.2.2 Expert Players
Expert human players performed significantly better than novices and Claude 3.5 Sonnet, with an average clustering score of 3.06 compared to Claude’s 1.78 (on the same 50 games). The distribution of weighted clustering scores is also far more right-skewed (see Appendix D.1 for more). Figure 6 shows that experts perfectly solved over 60% of the 50 games, while Claude 3.5 Sonnet fully solved 20% of these.
@@ 6 Discussion
6.1 What type of reasoning is hardest for LLMs?
To answer this question we rely on our taxonomy of reasoning types introduced in Section 3.2. The breakdowns of the reasoning types for the 1752 categories in our 438-game dataset are shown in Table 1. Figure 7 shows the performance of each LLM by reasoning type from our taxonomy of knowledge. Because the total counts of types of reasoning required across the 1,752 categories are unbalanced, the count of categories reasoned correctly is shown above each bar. The categories were only counted as correct if the model both clustered all the words correctly and justified its reasoning in a manner semantically analogous to the gold NYT Connections-provided category name.
The patterns in performance across the types of reasoning parallel the LLMs’ overall performances for the most part, with Llama 3.1 405B’s performance in Multiword Expressions and Word Meaning + Word Form (slightly better than GPT-4o) and Mistral Large 2’s performance in Word Meaning + Word Form (slightly better than Gemini 1.5 Pro) defying this pattern. The performance in reasoning categories across models is ranked from best to worst as follows: Semantic > Encyclopedic > Associative > Morphology/Orthography/Phonology > Multiword > Word Meaning + Word Form. This aligns with the game makers’ perceived level of category difficulty, as Word Meaning categories appear most often as yellow or green (easier) groupings, while Multiword Expressions and Word Meaning + Word Form categories are usually the purple groupings (most difficult). This is also consistent with the fact that the two types of reasoning LLMs performed best in—Semantic Relations and Encyclopedic Knowledge—exist in web-scraped information included in pre-training data, while other types of knowledge are more obscure and require iterative and deductive reasoning within each game. In every type of reasoning, however, LLMs perform with less than 50% accuracy. Though they may be clustering words together accurately, they are not always justifying these clusters correctly, and this is visible in the differences between the frequencies of unweighted clustering scores of 4 and categorical reasoning scores of 4 (Appendix D.2 Tables 5 and 6).
Since humans were not asked to provide justifications for their clusters, we do not include reasoning comparisons between LLMs and humans.
6.2 How do distractors prevent LLMs and humans from correct categorization?
The NYT Connections game is often formulated with item overlap in mind, according to the NYT Connections puzzle creator (Liu, 2023b). These distractors, or red herrings, make the game far more challenging. Red herrings can appear in two ways—as a red herring category or red herring word. In the former case, 3 ultimately unconnected words seem to form a category of their own with 1 word missing. In the latter, a category seems applicable to more than 4 words, but the extras belong to a separate grouping. Examples of each of these types of red herrings are in Appendix C.
@@ 6.3 How often do LLMs group the words correctly but present incorrect reasons?
| Gemini 1.5 Pro | 0.76 |
| Claude 3.5 Sonnet | 0.87 |
| GPT-4o | 0.84 |
| Llama 3.1 405B | 0.81 |
| Mistral Large 2 | 0.82 |
The average ratios in Table 2 are fairly high, close to or above 80%. The highest overall performing models Claude 3.5 Sonnet and GPT-4o have the highest ratios as well. Though it is fairly uncommon that a model will correctly group words without correctly naming the reasoning behind that grouping, there are very few instances where models received both a clustering score of 4 (fully solved game) and a categorical reasoning score of 4.
@@ 6.4 How can future work improve on such a benchmark?
Instead of choosing the first grouping, strategies grounded in System2 Thinking Evans (2003) could improve performance on such a benchmark. Generating multiple chains-of-thought reasoning and learning a model that assigns a higher reward to the correct reasoning chain prevents the model from arriving greedily to a suboptimal categorization. Allowing LLMs to solve the game one category at a time and incorporating the feedback present to humans in the original NYT Connections game, such as whether a grouping is correct (and what difficulty level it is by color), incorrect, or one word away from a correct grouping, may improve performance as well. Retrieval Augmentation from WordNet or dictionaries for lexical connotations Allaway and McKeown (2020) could further improve such categorization. Finally, creating synthetic training data and training an LLM on this task could further close the gap between expert human and LLM performance. We leave such exploration for future work.
@@ 7 Conclusions
We introduced the NYT Connections games as a benchmark to test abstract reasoning in state-of-the-art LLMs and evaluate their performance against expert and novice human players. We find that Claude 3.5 Sonnet performs best, although it is still no match for expert human players. By examining the performance through our knowledge taxonomy, we obtain a more solid understanding of areas in which LLMs can improve. Although most LLMs possess Word Meaning reasoning capabilities, they struggle with Multiword Expressions and combined knowledge categories. In addition, red herrings pose a challenge to current LLMs. Overall, we find that solving the NYT Connections requires a breadth of different knowledge types, which current LLMs do not seem to fully master.
@@ 8 Limitations
Many of the limitations in this section stem from the lack of data available forNYT Connections games and disparities in the comparison between LLMs and humans. Because it is a fairly recent invention and only one puzzle is released per day, there are only a few hundred games available. Since there are some category patterns learned through frequent play, ideally, a model trained on past NYT Connections games might bridge the gap between LLM and expert human evaluators’ performance.
We acknowledge that human evaluators were not required to add justifications for the groupings they made. This might have made performance comparisons between humans and LLMs for the types of reasoning more equal. Additionally, because a score of 3 was impossible in the human evaluation interface, we cannot be certain that humans were adept in the type of knowledge of their last category grouped, as this could simply have been a matter of grouping all options left. Other limitations of human evaluators include that because they were all peers or acquaintances of the paper’s authors, sampling bias could exist. Though the age range of the humans recruited was 14-60, other demographic factors that not have been accounted for in this sample.
@@ Appendix A Prompt
Categories will always be more specific than ‘5-LETTER-WORDS’, ‘NAMES’, or ‘VERBS.’
@@
Words: [‘DART’, ‘HEM’, ‘PLEAT’, ‘SEAM’, ‘CAN’, ‘CURE’, ‘DRY’, ‘FREEZE’, ‘BITE’, ‘EDGE’, ‘PUNCH’, ‘SPICE’, ‘CONDO’, ‘HAW’, ‘HERO’, ‘LOO’]
@@
Ways to preserve food: [‘CAN’, ‘CURE’, ‘DRY’, ‘FREEZE’]
@@
Sharp quality: [‘BITE’, ‘EDGE’, ‘PUNCH’, ‘SPICE’]
@@
Birds minus last letter: [‘CONDO’, ‘HAW’, ‘HERO’, ‘LOO’]
@@
Words: [1COLLECTIVE’, ‘COMMON’, ‘JOINT’, ‘MUTUAL’, ‘CLEAR’, ‘DRAIN’, ‘EMPTY’, ‘FLUSH’, ‘CIGARETTE’, ‘PENCIL’, ‘TICKET’, ‘TOE’, ‘AMERICAN’, ‘FEVER’, ‘LUCID’, ‘PIPE’]
@@
Shared: [‘COLLECTIVE’, ‘COMMON’, ‘JOINT’, ‘MUTUAL’]
@@
Rid of contents: [‘CLEAR’, ‘DRAIN’, ‘EMPTY’, ‘FLUSH’]
@@
Associated with “stub”: [‘CIGARETTE’, ‘PENCIL’, ‘TICKET’, ‘TOE’]
@@
__ Dream: [‘AMERICAN’, ‘FEVER’, ‘LUCID’, ‘PIPE’])
@@
Words: [‘HANGAR’, ‘RUNWAY’, ‘TARMAC’, ‘TERMINAL’, ‘ACTION’, ‘CLAIM’, ‘COMPLAINT’, ‘LAWSUIT’, ‘BEANBAG’, ‘CLUB’, ‘RING’, ‘TORCH’, ‘FOXGLOVE’, ‘GUMSHOE’, ‘TURNCOAT’, ‘WINDSOCK’]
@@
Parts of an airport: [‘HANGAR’, ‘RUNWAY’, ‘TARMAC’, ‘TERMINAL’]
@@
Things a juggler juggles: [‘BEANBAG’, ‘CLUB’, ‘RING’, ‘TORCH’]
@@
Words ending in clothing: [‘FOXGLOVE’, ‘GUMSHOE’, ‘TURNCOAT’, ‘WINDSOCK’]